OpenStudy (anonymous):

In a recent year, the total sales at Restaurant A were $75,000 more than at Restaurant B. Total sales at the two stores were 550,000,000. what were the sales for the year at each store. Restaurant B $________ Restaurant A $________

OpenStudy (anonymous):

okay you subtract 555,000 by 75000

OpenStudy (anonymous):

what did you get

OpenStudy (anonymous):

its 550,000,000 - 75,000= 549,925,000

OpenStudy (anonymous):

okay now it said that restaurant A 75000 dollars more than B

OpenStudy (anonymous):

Wait one sec

OpenStudy (anonymous):

is this multiple choice or fill in the blank

OpenStudy (anonymous):

feel in the blank

OpenStudy (anonymous):

In a recent year, the total sales at Restaurant A were $75,000 more than at Restaurant B. Total sales at the two stores were 550,000,000. what were the sales for the year at each store. Restaurant B $________ Simplify your answer. Type an integer or decimal Restaurant A $________ Simplify your answer. Type an integer or decimal

OpenStudy (anonymous):

okay lets see :) so: \[\text{Restaurant A = Restaurant B + 75,000}\]and\[\text{Restaurant A + Restaurant B = 550,000,000}\]so lets substitute the value of Restaurant A:\[\text{(Restaurant B + 75,000) + Restaurant B = 550,000,000}\]\[\text{2 \(\times\) Restaurant B + 75,000 = 550,000,000}\]subtract 75,000:\[\text{2 \(\times\) Restaurant B = 549,925,000}\]divide by two:\[\text{Restaurant B = 274,962,500}\]now plug in the value of "Restaurant B" into [Restaurant A + Restaurant B = 550,000,000":\[\text{Restaurant A + Restaurant B = 550,000,000}\]\[\text{Restaurant A + 274,962,500 = 550,000,000}\]subtract 274,962,500:\[\text{Restaurant A = 275,037,500}\] so that concludes that: \[\text{Restaurant A = 275,037,500}\]\[\text{Restaurant B = 274,962,500}\] Hope that helps! :) @MauriceLong

OpenStudy (anonymous):

I agree with @yummydum